OpenStudy (anonymous):

Multiple or divide as indicated. Simplify completely. (9z^3/y^5) / (8z^3y^5/5) This is what I've gotten, I was just wondering if someone could tell me if it was correct. I have a feeling it is not. I first switched the second numerator and denominator to change the problem into multiplication. I don't think anything is factorable. So, I then canceled out the z^3's but I'm not sure I can do that...Here's my work: 9z^3/y^5 * 5/8z^3y^5=48/8y^10 Thanks!

OpenStudy (johnweldon1993):

Perfect \[\large \frac{9z^3}{y^5} \times \frac{5}{8z^3y^5}\] \[\large \frac{45\cancel{z^3}}{8y^{10}\cancel{z^3}}\] \[\large \frac{45}{8y^{10}}\]

OpenStudy (johnweldon1993):

Your only (typo) was 9 times 5 = 45 not 48 as you put above

OpenStudy (anonymous):

Ok, sweet! Thank you so much @johnweldon1993 for taking the time to write all of that out. The help is much appreciated! :)